Assume that a procedure yields a binomial distribution with a trial repeated n=17n=17 times.

Use the binomial probability formula to find the probability of k=3k=3 successes given the probability p=p=1/4 of success on a single trial.

(Report answer accurate to 4 decimal places.)

P(X=3)=
Answer #1

Solution

Given that ,

p = 1 / 4 = 0.25

1 - p = 1 - 0.25 = 0.75

n = 17

x = 3

Using binomial probability formula ,

P(X = x) = ((n! / x! (n - x)!) * px * (1 - p)n - x

P(X = 3) = ((17! / 3! (17 - 3)!) * 0.2517 * (0.75)17 - 3

=  ((17! / 3! (14)!) * 0.2517 * (0.75)14

P(X = 3) = 0.1893

Probability = 0.1893